Use the Prison and Plea data in following table to calculate part a and b :...

  1. Use the Prison and Plea data in following table to calculate part a and b :

Guilty Plea

Plea of not Guilty

Sentenced to Prison

392

58

Not Sentenced to Prison

564

14

  1. If someone from the 1028 subjects is randomly selected, find the probability of selecting someone sentenced to prison.
  2. If someone from the 1028 subjects is randomly selected, find the probability of selecting someone sentenced to prison and entered a Guilty Plea.
Answer #1

a) A total of 1028 subjects

and total no of subject sentenced to prison=382+58=440

Hence the probability of selecting a subject sentenced to prison is

400.428

b. Total no of subject sentenced to prison is 440 and no of subject went for a guilty plea is 392+564=956

Hence Probability of selection subject sentenced to jail and went from a guilty plea is calculated as

= 0.381 1028